A fair die is rolled and a fair coin is tossed. The random variable X is defined as the number obtained on the die minus the number of heads (0 or 1) resulting on the coin. Calculate and write down a probability table for X.
Answer by robertb(5830)

The sample space consists of all ordered pairs (d,c), where d = 1,2,3,4,5,6, while c = 1,2.
The values of X = d-c are then 0, 1, 2, 3, 4, 5, and 6.
X = 0 appears once,
X = 1 appears twice,
X = 2 appears twice,
X = 3 appears twice,
X = 4 appears twice,
X = 5 appears twice,and
X = 6 appears once.
Since all the values on the die are equiprobable, and likewise with the coin, the probability table for X is as follows
X 0 1 2 3 4 5 6
p(X) 1/12 2/12=1/6 1/6 1/6 1/6 1/6 1/12